# Iterative Approach to Reconstructing Neural Disparity Fields from Light-field Data

### v1

@@ Abstract
This study proposes a neural disparity field (NDF) that establishes an implicit, continuous representation of scene disparity based on a neural field and an iterative approach to address the inverse problem of NDF reconstruction from light-field data. NDF enables seamless and precise characterization of disparity variations in three-dimensional scenes and can discretize disparity at any arbitrary resolution, overcoming the limitations of traditional disparity maps that are prone to sampling errors and interpolation inaccuracies. The proposed NDF network architecture utilizes hash encoding combined with multilayer perceptrons to capture detailed disparities in texture levels, thereby enhancing its ability to represent the geometric information of complex scenes. By leveraging the spatial-angular consistency inherent in light- field data, a differentiable forward model to generate a central view image from the light-field data is developed. Based on the forward model, an optimization scheme for the inverse problem of NDF reconstruction using differentiable propagation operators is established. Furthermore, an iterative solution method is adopted to reconstruct the NDF in the optimization scheme, which does not require training datasets and applies to light-field data captured by various acquisition methods. Experimental results demonstrate that high-quality NDF can be reconstructed from light-field data using the proposed method. High-resolution disparity can be effectively recovered by NDF, demonstrating its capability for the implicit, continuous representation of scene disparities.
@@
Neural Disparity Field, Disparity estimation, Light Field
@@ I Introduction
Computational imaging innovatively incorporates computer vision and signal processing techniques into imaging systems to reconstruct high-dimensional optical signals that carry scene information through novel imaging mechanisms, optical paths, and reconstruction methods, thereby achieving efficient and high-precision observation of scenes [1]. Computational imaging surpasses the limitations of classical imaging in terms of signal dimensions, scale, and resolution.
@@
From the perspective of imaging methodology, light-field imaging operates as a close-proximity passive direct line-of-sight imaging technology that relies on ambient light to capture scene content within its visual ambit. The characteristic features included narrower baselines, increased uniformity in viewpoint sampling, and diminished occlusions. Notably, substantial domain discrepancies emerge in the light-field data generated by different cameras owing to inherent variations in their imaging modalities, structural designs, and internal parameters.
@@
Deep-learning-based reconstruction methods rely on labeled training data and lack exploration of the inherent consistency within light-field data. Consequently, networks trained using data from one camera model may not be generalizable to other camera models. Given that individual light-field data contain abundant information, exploiting the full potential of their spatial-angular consistency empowers a network to reduce its reliance on specific camera models, thereby promoting generalization across different camera models.
The neural field [2] represents the scene to be reconstructed as a continuous implicit function, achieving advanced visual quality by the neural network parameterization. The continuous implicit function representation ensures continuity and is not limited by resolution constraints. By employing neural network parameterization of signal values, this representation method adopts a ”discrete-to-continuous” computational approach and utilizes differentiability for solving inverse problems. Compared with traditional methods, the neural field exhibits superior precision, lower complexity, and increased robustness. Furthermore, compared with deep learning approaches, their advantages include not requiring extensive training datasets, being unrestricted by data formats, having fewer parameters, and exhibiting high robustness.
@@
Therefore, we utilize a fully connected neural network to parameterize the disparity. The network constructs a neural disparity field (NDF) by incorporating hash encoding and multilayer perceptrons (MLPs) with activation functions. This approach implicitly represents the disparity contained in the light field and exploits the spatial-angular consistency within it. The inverse problem-solving model based on the NDF was established to reconstruct the disparity from the light field. This model was trained and iteratively solved through differentiability, effectively transforming the disparity estimation problem into an optimization problem for the NDF. The framework for the implicit representation of an NDF is shown in Fig.1.
@@
By exploiting the spatial-angular consistency within individual light-field data, the proposed method demonstrates exceptional performance in reconstructing the NDF. This reconstructed field can restore high-quality explicit discrete disparity maps. Unlike depth-learning-based methods for disparity estimation, the NDF reconstruction method does not depend on training datasets. It can be applied to two-plane parameterized light-field data using any collection method. By utilizing the implicit representation of the neural field, this disparity reconstruction method departs from the traditional approach of representing disparity as discrete points, overcoming the inevitable sampling, interpolation, and other discretization calculation issues of explicit discrete disparity and establishing a continuous disparity surface with implicit expression. The innovations of this study are as follows:
@@
    Based on the forward problem model, an optimization model for the inverse problem of the NDF was established using differentiable propagation operators. An iterative approach was proposed to solve the disparity reconstruction from the light-field. This method offers a novel solution to the disparity reconstruction problem. The iterative reconstruction method based on the inverse problem does not rely on training datasets and can be applied to two-plane parameterized light-field data obtained by any collection method. Furthermore, it allows the application of inverse problem-solving strategies in disparity reconstruction.
@@ II Related work
The 7D plenoptic function[3] was initially used to characterize the light fields. Over time, this function has been transformed into a two-plane four-parameter light field model[4], with the majority of commonly used methods relying on four-dimensional(4D) two-plane light fields[5]. The utilization of light field technology is crucial for providing depth information. To ensure generality, the proposed method is demonstrated using the central viewpoint depth and possesses the ability to reconstruct depth information from alternative viewpoints. Additionally, for light fields, there exists a one-to-one correspondence between the scene depth and disparity, which accurately represents the three-dimensional(3D) spatial structure of the scene. Existing methods can be broadly classified into two categories: rule-based traditional methods and deep-learning-based methods.
II-A Rule-based traditional methods
Disparity estimation algorithms that are rule-based traditional methods adhere to a computational approach known as ”discrete-to-discrete”. They aimed to estimate the corresponding discrete disparity images from discretized light-field data. However, this computational approach is influenced by the resolution limitations of disparity and unavoidably introduces various errors. A common source of error is sub-pixel interpolation, given that the accuracy of interpolation influences the estimation precision of the algorithm. For instance, Jeon et al. [6]proposed the use of the Fourier phase-shifting theory to replace bilinear interpolation, resulting in improved estimation accuracy. Currently, researchers have not noted advancements in the mitigation of discretization errors.
Early disparity estimation followed rule-based estimation methods, exploring the structural information of light-field data to assess the consistency among different views. Wanner et al. [7] introduced structural tensors into depth estimation from light-field data, enhancing the global consistency of the reconstructed depth by constraining the continuity of depth from different subview images of the light field. However, owing to the insufficient utilization of angular information in light-field data, the estimation performance of this algorithm requires further improvement. Heber et al. [8] introduced principal component analysis for the computation of stereo matching. They focused on subview images that impacted matching errors while disregarding those with minor influences. This approach improves the robustness of the algorithm and to some extent, its precision for stereo matching. To overcome the challenges of matching small baselines between subview images in light fields, Jeon et al. [6] applied Fourier phase-shifting theory to improve the accuracy of the matching computation. They adaptively aggregated the gradient costs based on the angular coordinates of the light fields and incorporated feature correspondences between subaperture images as additional constraints. Tao et al. [9]examined the properties of two depth cues derived from the correspondence and defocus in light-field data. Their findings demonstrate that these cues yield complementary depth information. Consequently, researchers have suggested merging the two depth cues to incorporate both angular and spatial information from the light field, thereby enhancing the performance of the algorithm for disparity estimation. However, this algorithm requires manual parameter adjustment to blend different results, thereby affecting its robustness across different scenes. In addition, it does not account for occlusion effects, leading to estimation errors in occluded regions.
To mitigate the errors arising from occlusions, Wang et al. [10] expanded the work of [9] by developing a model for the occlusion effect in light fields. They proposed an anti-occlusion algorithm by studying the relative positional relationships between different subview images in a light field and incorporating prior knowledge of object occlusion into stereo-matching algorithms. Although this algorithm partially addresses the challenges associated with occlusion in depth estimation by modeling occlusion effects in light fields, its straightforward occlusion detection criteria result in limited robustness across different scenes. To further mitigate the impact of occlusion on depth estimation performance, Sheng et al. [11] devised an estimation algorithm based on epipolar plane images (EPIs) combined with the local consistency of scene depth to address occlusion issues. Zhang et al. [12] considered the structure of EPIs and proposed a local depth estimation method based on a spinning parallelogram operator(SPO). Utilizing the regions segmented by the SPO in EPIs, they located lines representing depth information by maximizing the distribution distances between areas. Liu et al. [13] proposed an iterative scheme for scene depth reconstruction based on a 4D light field. The objective function of their optimization problem consisted of a fidelity term representing matching, along with penalty terms comprising gradient and classification terms. It iteratively corrects mismatched pixels to minimize the objective function, ultimately achieving accurate depth estimation. Subsequently, Liu et al. [14] mapped the pixel differences between matching regions to radial basis functions, where pixels with larger differences in occluded and noisy areas were mapped to smaller values in the radial basis functions, thereby effectively achieving anti-occlusion resistance and antinoise.
Although rule-based traditional disparity estimation algorithms achieve relatively high precision in depth estimation by exploiting the interrelationships among sub-aperture images, optimizing stereo matching computations, and addressing occlusion effects, their performance is still limited owing to the insufficient exploitation of angular and spatial information within the light field. Additionally, these methods, which use a ”discrete-to-discrete” approach, fail to overcome the resolution limitations inherent in the current representation of disparity and remain vulnerable to errors caused by data discretization. By contrast, neural fields with implicit representations, characterized by their inherent continuity, overcome the limitations imposed by finite-resolution discrete disparities and offer enhanced adaptability to complex variations in disparity and precision in their depiction.
@@ II-B Deep learning-based methods
With the advancements made in deep learning techniques for diverse computer vision tasks, there has been increasing interest in depth estimation from light-field data. Deep-learning-based algorithms primarily depend on supervised learning approaches, which require a large amount of training data annotated with accurate depth labels.
Heber et al. [15] introduced convolutional neural networks (CNNs) to address the task of depth estimation from light field data. Their approach utilizes supervised learning by providing a network with subaperture images in the horizontal and vertical directions and the corresponding depth labels. Data-augmentation techniques involving rotation, translation, brightness, and color adjustments have been proposed to improve the training process. Although the algorithm demonstrated superior performance on synthetic datasets compared to traditional algorithms, it required a large amount of depth-labeled training data, resulting in substantial data generation costs. In additionally, the domain gap between synthetic and real data poses challenges to the applicability of the algorithm in practical imaging systems. Subsequently, Shin et al. [16] proposed a multi-branch network structure with four branches dedicated to extracting features from light-field images in horizontal, vertical, and diagonally tilted directions. These features were then combined using CNNs to reconstruct the scene depth. Although the method showed improved estimation performance on synthetic data compared with traditional disparity estimation algorithms, it exhibited limitations in occluded and specular regions.
Alperovich et al. [17] developed a fully convolutional autoencoder for simultaneous depth and reflectance estimations. Although this supervised approach partially addresses the occlusion effects, its generalization ability is limited. Peng et al. [18] introduced unsupervised learning for depth estimation using light-field data. They proposed a combined loss function that imposed compliance and divergence constraints on the disparity between the warped sub-aperture images and the central view. This approach alleviates the challenge of acquiring labeled data and improves the generalizability of the algorithm across different imaging systems. Zhou et al. [19] utilized geometrically constrained EPIs to train networks and reconstruct depth information without requiring ground-truth depth labels. Li et al. [20] introduced an Oriented Relation Module that captured spatial relationships in vertical and horizontal EPIs. They enhanced training effectiveness and depth-estimation performance using refocusing-based data augmentation. Tsai et al. [21] presented a viewpoint-selection module that generated attention maps to determine the importance of each view for accurate depth estimation. This attention-driven approach improves the effective utilization of light field data. Li et al. [22] developed a lightweight model to estimate the disparity. The model employed physics-based multiscale cost-volume convolutional aggregation and an edge-guided subnetwork to enhance the geometric detail recovery near the edges and the overall performance. Huang et al. [23] designed an end-to-end trainable lightweight convolutional network that integrated cost volumes and attention modules. The designed network architecture enables the inference of light-field inputs at varying angles and facilitates detailed recovery in occluded regions.
The acquisition of the training and labeling data required by existing deep-learning algorithms poses a challenge, thus limiting their precision and applicability. Most deep learning-based approaches depend on data sourced exclusively from particular camera models during training, impeding their capacity to acclimate to various camera models and diminishing their generalization ability.
@@ II-C Neural fields
The neural radiance field [24] represents the scene to be reconstructed as a continuous implicit function parameterized by a neural network, achieving advanced visual quality and inspiring numerous subsequent works [25, 26, 27, 28, 29, 30]. Among these, mip-NeRF [25] reduces the multiscale representation of image aliasing in the neural radiance field. Mip-NeRF 360 [26] extended mip-NeRF to address poor performance in unbounded scene representations. Ref-NeRF [27] addresses specular and reflection issues in novel view synthesis. The instant-NGP [28] model enhances the training speed of neural radiance field models. TensorRF [29] decomposes the radiance field tensor into multiple lower-order tensor components to obtain a more compact scene representation. Factor Fields [30] is a unified framework based on coordinate-input networks, allowing the creation of powerful new signal representations and achieving improvements in approximation quality, compactness, and training time.
Continuous implicit functions parameterized by neural fields exhibit continuity and are not subject to limitations imposed by finite resolution, showing advantages over discrete representations (e.g., pixel color [31, 32], light field [33, 34, 35], signed distance [36], and occupancy [37]) in certain applications[38]. The core idea involves training MLPs with activation functions to map the input coordinates to the desired target values. Similar to existing coordinate-based representation methods, our method learns to map the coordinates of the central view to the corresponding disparities within the central view scene.

### v2

@@ Abstract
This study proposes a neural disparity field (NDF) that establishes an implicit, continuous representation of scene disparity based on a neural field and an iterative approach to address the inverse problem of NDF reconstruction from light-field (LF) data. NDF enables seamless and precise characterization of disparity variations in three-dimensional scenes and can discretize disparity at any arbitrary resolution, overcoming the limitations of traditional disparity maps that are prone to sampling errors and interpolation inaccuracies. The proposed NDF network architecture utilizes hash encoding combined with multilayer perceptrons (MLPs) to capture detailed disparities in texture levels, thereby enhancing its ability to represent the geometric information of complex scenes. By leveraging the spatial-angular consistency inherent in the LF data, a differentiable forward model to generate a central view image from the LF data is developed. Based on the forward model, an optimization scheme for the inverse problem of NDF reconstruction using differentiable propagation operators is established. Furthermore, an iterative solution method is adopted to reconstruct the NDF in the optimization scheme, which does not require training datasets and applies to LF data captured by various acquisition methods. Experimental results demonstrate that the proposed method can reconstruct high-quality NDF from LF data. The high-resolution disparity can be effectively recovered by NDF, demonstrating its capability for the implicit, continuous representation of scene disparities.
@@
Neural disparity field, Disparity estimation, Light-field
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ I Introduction
Computational imaging incorporates computer vision and signal processing techniques into imaging systems to reconstruct high-dimensional optical signals that carry scene information through novel imaging mechanisms, optical paths, and reconstruction methods, thereby achieving efficient and high-precision observation of scenes [1]. Computational imaging surpasses the limitations of classical imaging in terms of signal dimensions, scale, and resolution.
@@
From the perspective of imaging methodology, light-field (LF) imaging operates as a close-proximity passive direct line-of-sight imaging technology that relies on ambient light to capture scene content within its visual ambit. The characteristic features include narrower baselines, increased uniformity in viewpoint sampling, and diminished occlusions. Notably, substantial domain discrepancies emerge in the LF data generated by different cameras owing to inherent variations in their imaging modalities, structural designs, and internal parameters.
@@
Deep learning-based reconstruction methods rely on labeled training data and do not explore the inherent consistency within LF data. As a result, networks trained with data from one camera model are not generalizable to other camera models. Since individual LF data contain abundant information, exploiting their spatial-angular consistency allows a network to reduce its reliance on specific camera models, thereby promoting generalization across different camera models.
The neural field [2, 3, 4] represents the scene to be reconstructed as a continuous implicit function, achieving high-quality visual results through neural network parameterization. The continuous implicit function representation ensures continuity and is not limited by resolution constraints. By employing neural network parameterization of signal values, this representation method adopts a ‘discrete-to-continuous’ computational approach and utilizes differentiability for solving inverse problems. Compared to traditional methods like explicit pixel- or matrix-based representations, neural fields offer higher precision, lower computational complexity, and greater robustness. They achieve this by modeling scenes continuously, capturing details without resolution constraints. In contrast, traditional methods often suffer from aliasing or detail loss at higher resolutions. Neural fields are also more storage-efficient, utilizing implicit functions that reduce memory demands for high-resolution data. By focusing on the forward problem and using automatic differentiation to learn the inverse, neural fields adapt better to varying sampling patterns. Unlike deep learning methods, which require large labeled datasets, neural fields optimize individual scenes without such data. Their models are based on physical models rather than relying on ‘black-box’ mappings, allowing for improved generalization across datasets and better performance with new or unseen data.
The commonly used four-dimensional (4D) two-plane LF model[5, 6, 7] is shown in Fig. 1, where $(u,v)$ represents the viewpoint plane and $(x,y)$ represents the image plane. A pixel $(x^{\prime},y^{\prime})$ at viewpoint $(u_{i},v_{j})$ is expressed as $L_{u_{i},v_{j}}(x^{\prime},y^{\prime})$. A key feature of LF data is that all viewpoints lie on the same horizontal plane, with very small baselines between them. This small baseline makes three-dimensional (3D) reconstruction challenging, especially when using volume rendering methods[8, 2].
Figure 1: The four-dimensional two-plane LF model.
@@
Therefore, we leverage the spatial-angular consistency within individual LF data to implicitly represent the disparity contained within it and construct a neural disparity field (NDF). The inverse problem-solving model based on the NDF was established to reconstruct the disparity from the LF. This model was trained and iteratively solved through differentiability, effectively transforming the disparity estimation problem into an optimization problem for the NDF. The framework for the implicit representation of an NDF is shown in Fig.2.
@@
Unlike deep learning-based methods for disparity estimation, the NDF reconstruction method does not depend on training datasets. It can be applied to two-plane parameterized LF data using any collection method. By utilizing the implicit representation of the neural field, this disparity reconstruction method departs from the traditional approach of representing disparity as discrete points, overcoming the inevitable sampling, interpolation, and other discretization calculation issues of explicit discrete disparity and establishing a continuous disparity surface with implicit expression. The innovations of this study are as follows:
@@
    Based on the forward problem model, an optimization model for the inverse problem of the NDF was established using differentiable propagation operators. An iterative approach was proposed to solve the disparity reconstruction from the LF. This method offers a novel solution to the disparity reconstruction problem. The iterative reconstruction method based on the inverse problem does not rely on training datasets and can be applied to two-plane parameterized LF data obtained by any collection method. Furthermore, it allows the application of inverse problem-solving strategies in disparity reconstruction.
@@ II Related work
We review optimization-based and deep learning-based methods for LF depth estimation, as well as recent developments in neural fields.
II-A Optimization-based traditional methods
Disparity estimation algorithms based on traditional optimization methods adhere to a computational approach known as ‘discrete-to-discrete’. These algorithms aim to estimate corresponding discrete disparity images from discretized LF data. However, this approach is influenced by resolution limitations and inevitably introduces various errors.
Early disparity estimation methods explore the structural information of LF data to assess consistency among different views. Wanner et al. [9] introduce structural tensors for depth estimation. Heber et al. [10] employ principal component analysis for stereo matching, focusing on subview images that impact matching errors while neglecting those with minor influences. To overcome the challenges of matching small baselines between subview images in LFs, Jeon et al. [11] apply the Fourier phase-shifting theory to improve matching accuracy. Tao et al. [12] examine two depth cues derived from correspondence and defocus in LF data.
To address occlusion-related errors, Wang et al. [13] expand on [12] by developing a model that incorporates occlusion effects in LFs and propose an anti-occlusion algorithm that analyzes positional relationships between subviews and integrates occlusion knowledge into stereo matching. Sheng et al. [14] devise a depth estimation algorithm that combines epipolar plane images (EPIs) with local depth consistency to mitigate occlusion. Zhang et al. [15] consider the structure of EPIs and propose a local depth estimation method using a spinning parallelogram operator (SPO). By utilizing regions segmented by the SPO in EPIs, they maximize distribution distances to locate depth lines. Chen et al. [16] propose a depth estimation framework that regularizes the label confidence map and edge strength weights by detecting partially occluded boundary regions (POBR) and applying shrinkage/reinforcement operations. Liu et al. [17] introduce an iterative scheme for scene depth reconstruction based on a 4D LF, including a fidelity term for matching and penalty terms for gradient and classification components. Recently, Liu et al. [18] map pixel differences between matching regions to Gaussian functions, reducing the values of pixels with larger differences in occluded and noisy areas, thereby achieving anti-occlusion resistance and noise reduction.
In contrast to these traditional optimization methods, the proposed method employs implicit neural representations for continuous disparity modeling. This method effectively utilizes the spatial-angular consistency of LF data, overcoming the limitations of discrete data.
@@ II-B Deep learning-based methods
Advances in deep learning have increased attention to LF depth estimation, which can be categorized into supervised and unsupervised methods. Supervised methods rely on labeled depth to learn mappings from LF inputs to depth values, focusing on network architecture design. In contrast, unsupervised methods infer depth directly from LF data by leveraging its intrinsic structure and designing task-specific loss functions.
In supervised methods, early research typically predicted disparity information from LF data by designing feature extractors and employing ‘end-to-end’ network architectures. Heber et al. [19] introduced a CNN-based supervised method for depth estimation. Shin et al. [20] extracted features from LF data in various directions (horizontal, vertical, and diagonal) and combined them using CNNs for depth estimation. Alperovich et al. [21] developed a fully convolutional autoencoder for simultaneous depth and reflectance estimation. Li et al. [22] introduced an Oriented Relation Module that captures spatial relationships in EPIs and enhances training effectiveness through refocusing-based data augmentation. Tsai et al. [23] proposed a viewpoint-selection module that generates attention maps to assess the importance of each view, improving the utilization of LF data. Huang et al. [24] developed a lightweight convolutional network that integrates cost volumes and attention modules, enabling effective inference of LF inputs at varying angles and enhancing recovery in occluded regions. Li et al. [25] created a model for disparity estimation, employing physics-based multiscale cost-volume convolutional aggregation and an edge-guided subnetwork to enhance geometric detail recovery near edges and overall performance. Wang et al. [26] developed an occlusion-aware matching cost constructor that utilizes a series of convolutions with specific dilation rates, efficiently integrating pixels without shifting operations. At the same time, Wang et al. [27] proposed a generic mechanism to disentangle these coupled pieces of information for LF disparity estimation. Chao et al. [28] proposed a method for learning disparity distribution by constructing an interpolation-based cost volume at the sub-pixel level and designing an uncertainty-aware focal loss based on Jensen-Shannon divergence.
In unsupervised methods, Peng et al. [29] introduce a CNN-based approach for explicit depth estimation, using a combined loss function to align the disparity between warped sub-aperture images and the central view. Later, Peng et al. [30] propose a zero-shot learning-based framework for LF depth estimation. Iwatsuki et al. [31] redesign the loss function to indirectly measure disparity map accuracy by evaluating reprojection errors among the input LF views, and design pixel-wise weights and an edge loss to assess reprojection errors in the presence of occlusions. Zhou et al. [32] utilize geometrically constrained EPIs to train networks for depth reconstruction without requiring ground-truth depth labels. Furthermore, Zhou et al. [33] introduce an unsupervised LF depth estimation framework with explicit occlusion detection based on EPI geometry. Additionally, Li et al. [34] propose the Occlusion Pattern Aware Loss (OPAL), which successfully extracts and encodes the general occlusion patterns inherent in the LF for loss calculation.
While supervised learning relies on precise depth labels for training, unsupervised learning avoids the need for labels but requires complex loss functions and large datasets (except for zero-shot methods), along with extended training times. In contrast, the method proposed in this paper, based on implicit neural representation, operates without large-scale datasets or specific camera models, offering greater adaptability and efficiency.
@@ II-C Neural fields
The neural radiance field [2] represents the scene to be reconstructed as a continuous implicit function parameterized by a neural network, achieving high-quality visual quality and inspiring numerous subsequent works [35, 36, 37, 38, 39, 40]. Among these, mip-NeRF [35] reduces the multiscale representation of image aliasing in the neural radiance field. Mip-NeRF 360 [36] extended mip-NeRF to address poor performance in unbounded scene representations. Ref-NeRF [37] addresses specular and reflection issues in novel view synthesis. The instant-NGP [38] model enhances the training speed of neural radiance field models. TensorRF [39] decomposes the radiance field tensor into multiple lower-order tensor components to obtain a more compact scene representation. Factor Fields [40] is a unified framework based on coordinate-input networks, allowing the creation of powerful new signal representations and achieving improvements in approximation quality, compactness, and training time.
Chugunov et al. [41] tackle the challenge of obtaining high-fidelity depth maps in everyday photography by leveraging small baseline parallax from hand tremors, along with low-resolution LiDAR data. In a subsequent study [42], they propose an unsupervised approach that estimates high-quality depth and camera motion directly from 12-megapixel RAW frames and gyroscope data, relying on natural hand tremors for parallax cues. Although both methods aim to acquire depth from small baseline multiviews, they face challenges when applied to LF data. The minimal baselines between viewpoints and the equidistant sampling of the viewpoint plane result in the disparity between any two views being a multiple of the unit disparity relative to the viewpoint baseline. This constraint weakens effective disparity information and complicates disparity estimation using projection models. In the domain of Neural Disparity Refinement, Fabio et al. [43] propose a neural network-based method to refine noisy disparity maps generated by both traditional and deep learning stereo algorithms, aiming to improve depth estimation accuracy across various scenarios.
Our method’s core idea involves establishing a differentiable forward process that allows the neural network to learn, through backpropagation, to map the coordinates of the central view to the corresponding disparities within the scene. Unlike supervised or unsupervised methods, our approach is optimization-based, as it relies on an optimization process to iteratively refine the disparity estimation without requiring a training dataset.